OpenStudy (anonymous):

This is a Chain Rule problem: Evaluate the derivative of f(x= √sin2x at x= pi/4

OpenStudy (anonymous):

\[\sqrt{\sin(2x)} = (\sin(2x))^{1/2}\]

OpenStudy (anonymous):

I know that one, but do I plug in x after I find the derivative of the function or before finding it?

OpenStudy (anonymous):

after :)

OpenStudy (anonymous):

ahhhh, gotcha, thanks!